You have to wait for 3.5 years before starting to save for the purchase of a new car. You want to save towards the purchase of R12,500 car plus GST, Your savings rate is 5.2% compounded semi-annually and you want to be able to buy this car in 9 years. What is the size of your quarterly deposits?


1
Expert's answer
2021-10-12T09:38:04-0400
"FV=(1+i)P(\dfrac{(1+i)^n-1}{i})"

"P=FV\bigg(\dfrac{i}{(1+i)((1+i)^n-1)}\bigg)"

"P=12500\bigg(\dfrac{0.052/4}{(1+0.052/4)((1+0.052/4)^{(9-3.5)(4)}-1)}\bigg)"

"=488.11"

Your quarterly deposit is R488.11.